FEL Monthly Highlights
April 2001

On Saturday, April 21, the FEL team and approximately half of the Jefferson Lab staff educated (and
entertained) about 5,000 visitors from the local area during our 2001 Open House.  We had a continuous stream
of visitors from 10:00am to 4:00pm who were shown the FEL, exhibits in three of the user labs on basic and
applied applications of the FEL, and an exhibit on the planned addition of the Helios synchrotron to FEL Facility.

On the upgrade front, we now have ac successful model of our small dipoles that meet the field uniformity
requirements  (1/10,000).

JLab hosted a group of senior managers from Brookhaven National Laboratory to explore possible
collaborations in the areas of high current injectors and energy recovering linacs.  They are interested in exploring
the limits of these technologies.  We agreed to jointly pursue DOE, NSF, and NIH funding to support initiatives
in this area for application to cooling rings and light sources.  They are also interested in performing electron
transport experiments and code validation on the Upgrade as that becomes operational. They may be able to
provide some technical assistance in that effort.

A requested summary of our presentation to the Defense Science Board High Energy Laser Panel on Jan. 25th
was forwarded to the DSB support office this month.

We were visited this month by Dr. Al  Skolnick for a tour a and briefing on the FEL program. Dr. Skolnick was
involved with the management  of the Navy laser development program in the 1970-80's. George Neil and
Michelle Shinn were involved in numerous discussions this month  with Joint Technology Office (JTO) in the
preparation of the final FEL-materials exposure plan for the JTO study funded for this year.

F. Dylla attended the Spring 2001 Board of Trustees meeting of SURA and presented an update on the FEL
program.  We were pleased to have the Chief of Naval Research, Admiral Jay Cohen address the SURA Board
on Monday night.  The CNR discussed the potential interest of ONR in SURA's new initiative in coastal science
and gave mention to the existing ONR-SURA relationship for the development of FELs.  We have had three
discussions with the CNR since early February and during all three occasions he passed on his interest in
strengthening an FEL development program involving Jefferson Lab and potential partnerships with Navy  labs
and industry. George Neil gave presentations on the JLab FEL program this month at MIT, the Air Force
Institute of Technology in Dayton and at the Stanford FEL Laboratory. During the visit to AFIT, George had the
opportunity to brief managers in the AFRL in Dayton on our pending UV development program with the Air

The FEL program at Jefferson Lab was given high visibility at a Congressional briefing for the Virginia delegation
on April 26 which was organized by the Virginia Research and Technology Advisory Commission (VRTAC).
This is the third year that VRTAC has prioritized R&D activities in Virginia for the benefit of the delegation.  The
commission is co-chaired by Adm. Marc Pelaez, a former Chief of Naval Research and  Prof. Anita Jones
(UVA), the former assistant secretary of RDT&E for DOD.  Prior to VRTAC 's briefing on April 26, the current
CNR, Adm. Jay Cohen addressed the commission and again stressed his agency's interest in continuing FEL
development.  On Friday, April 27, we were visited by two current members of the Defense Science Board
panel on high energy lasers (Dr. Darrell Collier and Dr. Steve Hadley) for a briefing and tour of the FEL.

Project Cost Performance:
The project (Phase 1) budget for the period June 1, 2000 to September 30, 2001 is $9,029k.  The Phase 2
project was approved for $4,500k for a performance period of February 1, 2001 to September 30, 2002.  The
project through the month of April has a total of $3,461k of performance scheduled (assuming the project started
at the originally planned start date of April 1, 2000).  The work performed through the end of April was
$4,003k, which is 44% complete vs 38% scheduled.  The actual cost accrued through April totals $4,711k.
This results in a schedule variance of +$542k and a cost variance of -$708k.

WBS 3 (Beam Physics):
Initial performance analysis of the IR driver suggested the possibility of improved chromatic behavior. Minor
changes in sextupole placement led to much better orbit control off momentum.  Modeling of, and fitting on,
chromatic variations of propagated beam envelopes provided a significantly improved solution for matching
telescopes and produced a procedure for generating large acceptance matches. A broadband-chromatic model
was developed and applied to  optimize the energy-recovery endloop betatron matching.  The  resulting solution
provides space for Thomson scattering x-ray  generation apparatus. Field quality issues in the pi-bends and
injection line dipoles were under discussion.

A geometric quasi-perturbative model of the energy recovery arc is being generated to cross-check DIMAD
acceptance calculations.

WBS 4 (Injector):
Received the ion pumps and part of the flanges order for the gun chamber.  Arranged with the new W&M
student to start implantation work again. A test electrode will be implanted as warm-up for a Ti electrode which
will run first.

A review of the light box layout, revealed a correction allowing for a 2" aperture in the light box by going to a 4"
valve.  We  ordered a 100mm aperture RF shielded gate valve.  Drawings were completed for the cathode
shield. We will build and test the shield mechanism in UHV conditions.  Design changes for the ball cathode were
completed.  Ordered bellows for the stalk and shield modifications.  Details for the cathode support tube
modification have been completed.

WBS 5 (SRF):
Helium Vessels received from PHPK were qualified  and passed their QA test this month.  We have completed
the electron beam welding for two fundamental power couplers (FPC) bodies and qualified the procedure. A
kick-off meeting for the thermal shield vendor (ADC) was scheduled for April 25. Cavity vertical test assembly
(VTA) test began last week for the 7-cell cavity structures. Two tests have been completed after the standard
buffered chemical polish (BCP) treatment.  These cavities show onset of field emission at 11 MV/m. They will be
re-treated and re-tested until higher gradients are achieved.

WBS 6 (RF):
Zone 3 –
The HVPS is due 8/15/01.  The Circulators for this zone have been received and are being tested.  They will be
installed later this summer.
Zone 4 –
Door bracket was modified to close and latch properly.  The Cross-Connects for zones 2 & 4 are now wired
properly.  The wiring sheets were turned over to R. Vignato for documentation.  Marked-up drawings were
given to M. Gonzales for ECO's.  The Test Data book is being prepared.  Except for #8 circulator out for repair,
this zone is complete and waiting for a cryomodule.
Injector RF –
Factory Acceptance Tests of the 100 kW klystrons are still scheduled for 6/20/01.  The 100 kW Circulators will
be delivered to JLab early next week.  The only prospect, PSC, to upgrade the HVPS's for the 100 kW systems
is still waffling on his offer.  I am starting to make up a specification for the HV power transformer so that we will
be able to do this work.  Continued maintenance work on Quarter power supply.  Pursuing specifications for
transformer and inductors for Quarter.  The design drawings of the LCW manifolds for the klystron carts were
signed today..
Gun HVPS –
Glassman has received our spare unit and is working on its upgrade.  The large torus is a long lead item, so it will
several months before they will be finished and tested.

WBS 8 (I&C):
The "Alarms Handler" has been installed in the control room. A piece of software was added that looks at the
Beam Loss Monitors (BLMs) and prompts the operator if an attempt is made to open the drive laser shutter
while the machine is in Beam Mode 3, or 4 (unsafe condition). These updates will help the transition of
operations from the FEL core group to "fresh" operators. (Not "inexperienced" since they have a great deal of
experience with CEBAF)  The server hard drives were also upgraded.

Design activity for a VME based Beam Loss Monitor (BLM) card/system is ramping up. Informal discussions
are ongoing regarding addition features. The Integrator "front-end" electronic will remain essentially untouched
but an additional integrator will be added for a "low beam loss" diagnostic. This second integrator will be
controlled by a DSP, the desire is to have a loss calculated and displayed in units something like
"nanoamp-seconds". The design of the new four channel VME timing/delay module continues, the design of the
MPS Input card is proceeding as well.

Initial planning has gone into interfacing the YAG laser in Lab 2 for the Optics group.  The vacuum system layout
Rev. 0 is complete. Drawing are being prepared then they will be reviewed. The process of gathering parts for
the Ion Pump Power Supplies is nearly complete. Received bids from Allied and Newark on parts that are still

Bid requests have been sent out to 5 Contractors for assembly of the User power supplies for the labs. This will
also help to establish a good source for other up-coming fabrication jobs. The remaining VME front panels that
were sent out for machining and silkscreen are back. All spare parts for the VIDMAR in Lab 5
have been ordered. The autotransformers arrived, these are to act as a back-up in case of a failure of the
50KWatt HPA variacs. These were responsible for a number of hours of down time during the last run.

EECAD drawings for the Charge Totalizer Current Monitor and the LSS Alignment Mode Protection Card were
signed off as complete.  The Charge Dump Current Monitor Chassis Fabrication, Assembly and Wiring
Diagrams are out to be checked.  The picomotor controller chassis assembly and fabrication drawings are also
Cable tray has already been ordered. LCW and vacuum layouts/requirements are also progressing well.
The Laser Safety System (LSS) master has been updated for operating multiple hutches in any or all of the user
labs. There is now available status for operation of up to 5 hutches in the six labs. An "FEL beam-block" shutter
has been developed by the optics group that will allow each hutch to be manually bypassed for rapid access and
sample changes without dropping the labs LSS.  This beam-block and the user lab shutter provide the two
independent methods of preventing severe biological damage from the FEL beam to the user. The Lab 2 LSS is
installed and complete. Final certifications for all of the LSS components require the FEL to be operating so this
will be done just prior to the summer run. (Note: preliminary certs have been done as a first check and as a
training exercise).

Parts for a prototype BPM using Motorola DSP56F807 and serial CAN bus including filters and
down-convertors have been ordered. The software development kit is in for this DSP. The BPMs are a couple
weeks late from Modern Machine but do not yet threaten the schedule. 20 feet of table space has been setup in
the back of the FEL for the electrical assembly of the beam viewers prior to passing them off for "girder"
installation. During the last go-round this was done in the tunnel- we're learning.  The new design for the OBPM
PCB using the Molectron detectors is complete and assembled, and will be tested.

First 12 Beam Position Monitors have been delivered. 2 of the viewers are also here, Master Machine assures us
that they will catch-up. There is no impact on the schedule other than my desire to get these behind us, girder
assemblies will not start for ~ 3 months. A clean hood will be added to Area 51 in the FEL to allow for dust free
assembly of the "odds & ends". Some of these are Michelle's Ultra-Viewers!, and the wiggler viewers. James
continued to spend most of the week finishing up on the LSS system upgrades.  The new box in Lab 2 has been
installed and tested; the LSS Master Chassis has been upgraded to include all six user labs; plans are in progress
for adding additional protective beam stops to lab 3 for multiple hutches.  The hutch in lab 6 is being equipped
with multiple magnetic switches to ensure all doors on the hutch are shut prior to beam operation in that lab when
in the hutch mode.

WBS 9 (Transport):
Injector Dipoles (DU/DV)
o DULY Research used their experience with their RADIA magnetic analysis code, their newly upgraded, high
    memory computers and a new element generation algorithm to investigate the field flatness and integral flatness on
    the large injector dipole (GU).  They achieved success, obtaining values for both characteristics of less than 1
    part in 10,000 over the good field region.  In addition, the noise in their analysis, the mathematically generated
    oscillations were all below these values.
o DULY Research was finally successful in getting their magnetic model of the small injector Dipole (GV) to
    achieve the required tolerances of 1 part in 10,000 for both field flatness and field integral.  Their trick for this
    “all-ends” magnet was to add a set of trim coils that raised the saturation of the iron to achieve field uniformity
    and to have magnetic shorts on the field clamp to achieve field integral uniformity.
o They are working on incorporating the findings of the magnetic model of the GV into a practical hardware
o They have completed drafts of the yoke details of the GU. They continue work on the details of the coils.
Optical Chicane Dipole (DW)
o We re-checked the drawing set and the back checks were incorporated.  .
o We bought sample strips from five rolls of electrical grade steel that is destined to be used as the pole tips on
    these optical chicane dipoles as well as the 180° dipoles.  We measured their thickness over the very flat center
    of the sheets, selected the best roll and bought enough sheets for all our magnets.  (They have to be flat to tenths
    of thousandth of an inch).
o Leigh Harwood checked the specification for the major dipole steel purchase, it was signed off and the
    package went out for bid.
o A contract was signed with Innovation Services, a consulting firm specializing in electrical insulation systems
    using epoxy, to review our dipole coil specification and bring it up to date for the topics of epoxy formulation,
    curing, manufacturing and quality control.
Arc Dipoles (GY, GX, GQ)
o Advanced Energy Systems (AES) almost finished designing the 180 degree dipole (GY). They laid out the
    leads, hoses, cover, manifolds and terminal strips for the path length correctors. At months end they stopped
    further detailing pending the final magnetic model findings.
o A review of the diagnostics requirements led to a new requirement.  Two sets of holes through the yoke plates,
    near the coils at around 54 degrees from the face will be required. These holes will accommodate two periscopes
    that look at each other through windows in the outside surfaces of the vacuum chamber.  The two periscopes will
    allow a HeNe Laser to establish known optical path that coincides with the synchrotron light generated at the
    apex of the magnet when the beam is centered.  The holes through the yoke plates are similar to the single set
    used in the IR Demo.  That set produced no measurable magnetic anomaly.
o The magnetic analysis of the GY with path length correctors ran into a snag as Dave Douglas found that
    increasing the bulk field to correct for the lower field in the path length corrector slot does not meet the field
    tolerance of 1 part in 1000 imposed by dispersed beam. Tom Schultheiss is calculating how much additional face
    steel is needed and how much parallel offset there is to the beam if it is brought into the magnet perpendicular but
    off set and comes out in a similar condition.  David says he can live with a small amount  (say 1.5 mm) of off set.
    Preliminary results however yield 3 mm.  The final solution may be a compromise between increasing the bulk
    field somewhat and allowing a small off set.
o They have begun design and magnetic modeling the Arc Bend Dipole (GX) after some go-rounds on the exact
QG (3 inch quad)
o The coils manufacturing specification was checked, coil drawings signed and the coil package sent out to
    vendors for bid.
o The prototype core was sent to the shop for a first pass at determining the pole end chamfer needed to cancel
    out dodecupole
 QX (Trim Quad)
o A vendor has agreed to custom make conductor for these quadrupoles to a defined dimension to a very high
    tolerance.  This should considerably decrease the manufacturing problems associated with conductor size that
    have plagued us in the past.  This is necessary because tolerance build-up in this multi-turn, conductor geometry
    dominated magnet will ruin the field gradient quality.
o We created a working magnetic model and are now debugging its meshing.
Corrector Dipoles (DB & DJ)
o After a low bid was obtained from a new vendor.  However, we discovered that the drawings on the bid were
    sent out without a needed specification.  A new package, (that will be easier for vendor’s to understand) is in
    vendor’s hands for are-bid.
o We continued laying out the girders in 3D IDEAs format and started laying out the girders and stands, making
    the tradeoffs of machined-in features vs. adjustment mechanisms
o We designed more of the beam chamber in the arcs, firming up the parameters associated with chamber height.

o We sent out inquires to vacuum vendors to determine if they can make our style of chamber and received many
    responses.  The JLab Shop, which did much of the chamber work last time, may be overloaded at our required
    time of fabrication.
o We took our first look at piping the LCW in the upgrade.
o The alignment group at CERN is going to sell us 6 of their heavy weight rated alignment cartridges for use
    under our 180° dipoles so we will not have to go through the engineering.

WBS 10 (Wiggler):
A preliminary set of magnet measurements were performed on the dispersion section and verified the first order
field requirements.  The magnet must be disassembled to add holes that the vendor missed on the support stand.
Detailed magnet measurements will start once the magnet is reassembled.

Estimates were received on the magnetic modeling of the wigglers.  The estimates were very high so it was
decided to figure out the strength of the end coils by direct measurements.  We decided to analyze the end coils
of the Wiggler ourselves because the cost to have AES do it was much greater than simply winding conductor on
the poles and deducing the proper number of turns from measurement (Analog computer method).

A coil design for the wiggler end correctors has been submitted to the shop.  In the meantime we will be taking
measurements of the DC field in the wiggler vs. the strength of the end correction using some AT coils left over
from CEBAF.  This should allow us to determine the excitation necessary for these coils.  All components for the
wiggler viewers have now been received.  The small tubing for the vacuum chamber has also been received.  We
should be able to start machining of the chamber soon.

WBS 11 (Optics):
Meetings were held early in the month to discuss mirror mounting concepts for the optical cavity assemblies.  The
group concurred with our scheme for employing 4 mirrors mounted in JLab-designed mounts, and so we're now
designing them.  An internal document was distributed for comment that details the design evolution of the optical
cavity.  The ultraviewer (alignment) design for 4" tubing is being detailed. Procurement received three bids to
build two prototype deformable mirrors; we anticipate an award early this month. A meeting was held to discuss
how best to accommodate ultrashort x-ray production and delivery in the Upgrade, and still accommodate
optical and electron-beam diagnostics downstream of the wiggler. The impact for optical diagnostics was that we
will no longer plan on fastening our diagnostic onto the wiggler vacuum chamber, but will instead design it on a

The 50% mirror in the User Lab 1 mirror cassette was realigned to make it's beam coincident with the other two.
Other vacuum parts needed for the User Lab 2 mirror cassette were received. I & C, in conjunction with the
electrical installation group, wired smoke detectors in the drive laser enclosure, and in User Lab 2.
Other LSS work also occurred.

The optical cavity assembly conceptual design was reviewed and discussed; several good points were raised
regarding internally supporting the linear translation, and the stiffness of the mirror mounts.  These comments are
being incorporated in the design. The ultraviewer (alignment) design for 4" tubing is being detailed. We also
reviewed it and came up with an action item to look at an alternate baseplate fabrications scheme.  We agreed on
a particular mirror orientation (out of two possible).  We also agreed to add an extra cross (with no components)
next to the ultraviewers, to allow for other optical components/diagnostics to be added at a future date.
Top-level specs on the diagnostic beam dump for the FEL were distributed, and the requisite make vs buy
comparison is being made.  Procurement awarded the contract to build two prototype deformable mirrors.

Other parts for the manually-actuated safety shutter arrived and were fitted, the design works well.  A few
modifications to hold the limit switches need to be made, and then we can try it in User Lab 3 The LSOP for
User Lab 2 was submitted for sign off.

The optical cavity assembly design progresses, it now has a stiffer support frame and accommodates an exterior
motor for translation.  The check prints for the new ultraviewer will be produced today.  The manually-actuated
turning mirror (initially for the User Lab 3) was submitted to our machine shop for fabrication.  I & C was
supported in their efforts to test out the LSS.  The LSOP for User Lab 2 has been signed off within the FEL
Dept., and is now with Safety Dept. officialsWe had Robert Chow (LLNL) here on Tuesday to evaluate our
plans for the Mirror Test Stand (MTS).

Overall, he thought we had a very good plan.  He did alert us to some experimental subtleties, and offered to
share some designs that proved successful in the high average power laser system (AVLIS) operated at LLNL in
the '80s & '90s.  We anticipate receipt of the rest of the MTS internal hardware next week.  We reviewed the
check prints for the new ultraviewer, and identified several changes; these are being incorporated.  We should
proceed to signoff next week.  Conceptual designs for the diagnostic beam dump and the collimator proceeds,
for the latter, S. Benson has done some optical modeling that defines the types of mirrors we'll need.

Parts for the manually-actuated turning mirror (initially for the User Lab 3) were received from our machine
shop.  Assembling them identified a few minor interferences, these are being fixed.  Supports for the beamline are
being designed and procured.  Some work was done on the drive laser to realign it's output; it will be needed
next week for OBPM tests.  The LSOP for User Lab 2 has been signed off.  After we receive some beamline
shields, we'll begin testing the laser, and then begin our first checkouts of the MTS.

A training meeting was held on the optical systems with the MCC operations coordinator to explain functionality
and characteristics of the cavity and optical transport system.
FEL X-ray Research:

    1. Boyce prepared an LOP for the next run. There are two objectives: a) measure spatial distribution and b)
        investigate the anamolous x-ray peak found in the last run.
    2. In order to accomplish the above, some equipment needs to be modified and I have designed a Pb shield
        for the X-ray detector. I've scrounged two Pb bricks that have never been exposed to radiation and supplied the
        shop with them along with the design. Casey anticipates the job to be complete by  next Wednesday.
    3. In parallel with the shop activities, we're working to modify the detector translator station that will move the
        x-ray detector across the Be window through which the diffracted x-rays emerge.
    4. Met with Steve Benson, Michelle Shinn, and Dave Douglas to discuss the X-ray requirements for the
        upgrade. We have decided to limit the capabilities to x-rays below 50 keV. Further, the quad triplet can now be
        moved closer to the wiggler and the x-ray chamber will be between that quad triplet and the next quad triplet. I
        am working on a design of the chamber.

Planning continues for the up-coming runs to support the JTO proposal and a call for proposals for the summer
run is being prepared. We heard from Peter Schaaf (Univ. of Goettingen) that the metal nitrides samples that he
made during the recent winter run appear to be very high quality nitrides.